### v1

# Thermodynamic Signatures of Gaussian Entanglement Beyond Entropy

Abstract
Continuous-variable quantum thermodynamics in the Gaussian regime provides a promising framework for investigating the energetic role of quantum correlations, particularly in optical systems. In this work, we introduce an entropy-free criterion for entanglement detection in bipartite Gaussian states, rooted in a distinct thermodynamic quantity: ergotropy—the maximum extractable work via unitary operations. By defining the relative ergotropic gap, which quantifies the disparity between global and local ergotropy, we derive two independent analytical bounds that distinguish entangled from separable states. We show that for a broad class of quantum states, the bounds coincide, making the criterion both necessary and sufficient. We further extend our analysis to certain non-Gaussian states and observe analogous energy-based signatures of quantum correlations. These findings establish a direct operational link between entanglement and energy storage, offering an experimentally accessible approach to entanglement detection in continuous-variable optical platforms.
Introduction.– Quantum thermodynamics examines how intrinsically quantum resources shape energy and information processing in microscopic regimes [1, 2, 3, 4, 5, 6, 7, 8, 9]. With experimental advances enabling precise control over single quantum systems, addressing the energetics of nanoscale devices has become both timely and essential [10]. A key concept in this context is ergotropy— the maximum amount of energy that can be extracted from a quantum system through unitary evolution [11]. The importance of this quantity lies in its sole dependency on the Hamiltonian or energy, which essentially captures a distinct, entropy-free notion in the context of quantum resources [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
Continuous-variable (CV) systems, particularly quantum optical platforms, offer a compelling setting to explore these thermodynamic aspects. CV states are central to diverse quantum technologies, including quantum communication, sensing, and computation [28, 29, 30]. Among these, Gaussian states are of particular importance due to their experimental feasibility and the efficient implementation of Gaussian operations [31]. These features make them an attractive testbed for probing the energetic role of quantum correlations. Entanglement in Gaussian systems is typically assessed using entropic measures [32] or criteria such as the positive partial transpose (PPT) condition [33]. However, such approaches are grounded in information-theoretic or purely mathematical constructs and do not directly connect to physically measurable energy quantities.
This motivates the development of energy-based entanglement witnesses that operate independently of abstract informational and mathematical metrics. Recent work suggests that ergotropy offers a thermodynamically meaningful framework for revealing quantum correlations [16, 17, 18, 19, 21, 22, 23, 24, 25, 26, 27, 20]. However, these ideas have primarily been explored in discrete-variable (DV) systems. Extending them to CV systems is both conceptually relevant and practically valuable, particularly in the context of emerging quantum technologies, such as quantum batteries, where Gaussian optical systems could provide a scalable and controllable platform [34].
In this work, we introduce the relative ergotropic gap (REG), a quantity that captures the difference in extractable energy between a global bipartite Gaussian state and its locally accessible marginals under Gaussian unitary evolution. We derive analytical bounds that distinguish entangled from separable states, and demonstrate that for a broad class of Gaussian states, REG serves as a necessary and sufficient criterion for entanglement. Operationally, this reveals that entangled states can store and release energy in ways fundamentally beyond the reach of separable states. We further extend our analysis to selected non-Gaussian regimes [35], namely photon-subtracted states, demonstrating that large REG implies higher entanglement (as quantified by the Shchukin-Vogel criterion [36]). These results point toward a broader applicability of energy-based signatures for entanglement in CV systems, with direct relevance to the development of energy-efficient quantum technologies.
Energy Extraction.– In thermodynamics, the amount of work extractable from a system is controlled by its internal energy and entropy. For adiabatic processes, work extraction depends solely on internal energy changes. In the quantum regime, the analogous process is unitary evolution, an isolated dynamics that preserves von Neumann entropy. Ergotropy is the maximum amount of work reversibly extractable [37, 38, 11, 39]. For a state $\rho$ with Hamiltonian $H$, it is defined as $\mathcal{E}(\rho):=E(\rho)-\min_{U}E(U\rho U^{\dagger}),$ where the minimization is over all unitaries $U$, and $E(X):=\operatorname{Tr}(XH)$ denotes the system’s energy in state $X$. This expression quantifies the energy difference between $\rho$ and its passive counterpart $\rho^{p}$, the state with the same spectrum but minimal energy, beyond which no work can be extracted through unitaries. Ergotropy thus quantifies how far a state is from being passive, in purely energetic terms and without reference to entropy.
Energy Extraction in CV Systems.– In systems with infinite-dimensional Hilbert spaces, the implementation of a generic unitary is practically challenging. However, restricting to Gaussian unitaries leads to the concept of Gaussian ergotropy [40, 41], defined as
$\mathcal{E}_{G}(\rho):=E(\rho)-\min_{U_{G}}E(U_{G}\rho U_{G}^{\dagger})=E(\rho)-E(\rho^{p}_{G})$ (1)
where $U_{G}$ denotes a Gaussian unitary and $\rho^{p}_{G}$ is the corresponding Gaussian passive state that minimizes the system’s energy under such operations. It is evident that for an arbitrary (possibly non-Gaussian) state $\rho$, Gaussian ergotropy satisfies $\mathcal{E}_{G}(\rho)\leq\mathcal{E}(\rho)$, whereas if $\rho$ is itself a Gaussian state, this inequality becomes an equality because Gaussian unitaries suffice to reach the true passive configuration. In such cases, we omit the subscript G and simply refer to ergotropy.
In composite or multi-mode CV systems, Gaussian ergotropy can be defined globally (via joint Gaussian unitaries $U_{G}$) or locally (via product unitaries $\otimes_{i}U_{i}$, where $U_{i}$ acts on $i^{th}$ mode). Their difference defines the Gaussian ergotropic gap:
$\Delta\mathcal{E}_{G}$ $:=\mathcal{E}^{\text{global}}_{G}(\rho)-\mathcal{E}^{\text{local}}_{G}(\rho)=E_{lp}-E_{gp}$ (2)
where $E_{lp}$ and $E_{gp}$ denote the energies of the local and global Gaussian passive states, which are obtained via separable local unitaries $\otimes_{i}U_{i}$ and global Gaussian unitaries $U_{G}$, respectively. Operationally, this quantity captures the energetic advantage of performing correlated (global) operations over local ones. While the notion of ergotropic gap has previously been explored in DV settings as a thermodynamic signature of multipartite quantum correlations [17, 18, 19, 22], its extension to CV systems remains largely unexplored.
Gaussian states constitute the backbone of optical quantum technologies and offer the operational advantage of being completely characterized by their first and second statistical moments. Their simplified mathematical structure enables analytic treatment of thermodynamic quantities such as energy, entropy, and ergotropy, making them a natural playground for quantum thermodynamics in CV systems [29, 40, 41, 42].
In phase space, an $N$-mode bosonic system is described by the quadrature operators $\hat{\vec{r}}=(\hat{x}_{1},\hat{p}_{1},\ldots,\hat{x}_{N},\hat{p}_{N})^{T}$, which satisfy the canonical commutation relations $[\hat{r}_{j},\hat{r}_{k}]=i\Omega_{jk}$, where $\Omega=\bigoplus^{N}i\sigma_{Y}$ is the symplectic form [29] and $\sigma_{Y}$ is the Pauli $Y$ matrix.
The energy of a Gaussian state depends on its covariance matrix $\sigma$, with entries $\sigma_{jk}=\langle\{\hat{r}_{j},\hat{r}_{k}\}\rangle-2\langle\hat{r}_{j}\rangle\langle\hat{r}_{k}\rangle$ and the system’s Hamiltonian, which, for a system of quadratically coupled harmonic oscillators with frequencies $\omega_{k}$, can always be brought to the form $\hat{H}=\sum_{k}\omega_{k}(\hat{x}_{k}^{2}+\hat{p}_{k}^{2})$. The average energy of a Gaussian state $\rho$ with covariance matrix $\sigma$ is then
$E(\rho)=\frac{1}{4}\sum_{k}\omega_{k}\left(\text{Tr}[\sigma_{k}]-2\right),$ (3)
where $\sigma_{k}$ is the $2\times 2$ covariance submatrix corresponding to mode $k$. Note that first moments also contribute to the energy of a Gaussian state, but since they can be removed through local unitary operations, they do not affect the ergotropic gap — which quantifies the energy difference between local and global passive states — and thus can be disregarded in our analysis. For a two-mode Gaussian state, one can always bring the covariance matrix to the so-called standard form $\sigma_{sf}$ via local Gaussian unitary operations:
$\sigma_{sf}$ $=\begin{pmatrix}a&0&c_{1}&0\\ 0&a&0&c_{2}\\ c_{1}&0&b&0\\ 0&c_{2}&0&b\end{pmatrix},$ with $\text{with }a\geq b\geq 1\text{ and }c_{1}\geq|c_{2}|.$ (4)
This form is particularly useful for computing both local and global passive energies. The covariance matrix $\sigma_{lp}$ of the local passive state is obtained via local Gaussian operations, while the global passive covariance matrix $\sigma_{gp}$ is achieved via a symplectic transformation that diagonalizes $\sigma_{sf}$ into its Williamson form: $\sigma^{gp}=\text{diag}(\nu_{+},\nu_{+},\nu_{-},\nu_{-})$, where $\nu_{\pm}\geq 1$ are the symplectic eigenvalues of $\sigma_{sf}$, i.e. the moduli of the eigenvalues of $i\Omega\sigma_{sf}$. Assuming $\omega_{A}\leq\omega_{B}$, the ergotropic gap for a two-mode Gaussian state $\rho_{G}$ reads:
$\Delta\mathcal{E}:=E_{lp}-E_{gp}=\frac{1}{2}[\omega_{A}(a-\nu_{+})+\omega_{B}(b-\nu_{-})]$ (5)
where these local and global passive energies are computed directly from the covariance matrices, since they fully determine the passive state’s properties in Gaussian systems.
Main Results: Ergotropic Gap versus Correlations of Gaussian States.– We begin by examining the interplay between quantum correlations and work extraction in the case of pure two-mode Gaussian states. This regime is particularly amenable to analytical treatment and provides a well-understood setting, as the entanglement of pure bipartite states is fully characterized by the entropy of the reduced state.
Theorem 1.
For pure two-mode Gaussian states $\rho$, the ergotropic gap $\Delta\mathcal{E}(\rho)$ vanishes if and only if $\rho$ is separable. Moreover, $\Delta\mathcal{E}(\rho)$ is a strictly increasing function of the mutual information $I(\rho)$, and hence provides an equivalent quantification of entanglement for these states.
The proof is discussed in Appendix A. Similar to the DV case [18, 19], we find that the ergotropic gap is as a valid entanglement measure for pure two-mode Gaussian states. Since these states are fully characterized by a single parameter, all entanglement measures must be mutually dependent and equivalently capture the quantum correlations. This reflects the situation in pure $2\times d$ dimensional DV systems, where all entanglement measures can be related to the entanglement entropy [43, 44]. The result in Theorem 1 readily extends to pure bipartite Gaussian states of an arbitrary number of modes, by virtue of their mode-wise decomposition into products of two-mode states via local unitaries [45].
The natural question arises regarding whether the ergotropic gap could also serve as a reliable entanglement quantifier for mixed states. Before addressing this, we examine whether $\Delta\mathcal{E}(\rho)$, viewed as a mathematical functional, satisfies basic criteria for a correlation measure [46]. In DV systems, the ergotropic gap has been shown to witness entanglement, but it can be nonzero even for product states [18]. Interestingly, our next result reveals a distinct behavior in the Gaussian setting:
Lemma 1.
For general two-mode Gaussian states, $\Delta\mathcal{E}(\rho)>0$ holds if and only if the state exhibits correlations.
The proof is provided in Appendix D. This establishes that $\Delta\mathcal{E}(\rho)$ acts as a faithful indicator of correlations, being strictly positive if and only if the state is non-product. We now turn to the case of correlated mixed states, exploring whether this quantity can consistently reflect the strength of quantum correlations.
As a preliminary case study, we consider the well-known two-mode squeezed thermal state (TMS), obtained by applying a two-mode squeezing operation with parameter $r$ to two bosonic modes of frequency $\omega$, each initially in thermal equilibrium with a heat bath at temperature $T$. The resulting state’s covariance matrix is itself in standard form:
$\small\sigma_{\text{TMS}}=\left[\begin{matrix}k\cosh{\left(2r\right)}\mathds{1}&k\sinh{\left(2r\right)}\sigma_{z}\\ k\sinh{\left(2r\right)}\sigma_{z}&k\cosh{\left(2r\right)}\mathds{1}\end{matrix}\right],\,\,\text{where}\ \ k=\coth\left(\mbox{$\frac{\omega}{2T}$}\right)\!.$
Its corresponding gap is $\Delta\mathcal{E}_{\text{TMS}}=2k\omega\sinh^{2}r$, which grows monotonically with temperature and diverges as $T\to\infty$. This behavior is particularly striking because thermal noise is generally expected to degrade quantum correlations. Indeed, the quantum mutual information decreases with increasing temperature and remains bounded in the limit of infinite thermal noise: $\lim_{k\rightarrow\infty}I(\sigma_{\text{TMS}})=2\log{\left(\cosh{\left(2r\right)}\right)}$.
This divergence of $\Delta\mathcal{E}$ underscores a fundamental distinction between CV and DV systems. In DV systems, rising temperature leads the state toward a maximally mixed configuration—a global passive state—causing the ergotropic gap to vanish. In contrast, in CV systems, the unbounded energy spectrum allows the thermal population of arbitrarily high energy levels without necessarily rendering the state passive. Consequently, even minimal correlations in extremely energetic states can lead to diverging ergotropic gaps despite diminishing mutual information.
This makes the ergotropic gap $\Delta\mathcal{E}$, in its vanilla form, an unfit candidate to quantify correlations in CV mixed states. To mitigate the divergence problem, we introduce the relative ergotropic gap (REG), defined as
$\Delta\mathcal{E}_{\text{rel}}\equiv\frac{\Delta\mathcal{E}}{E_{gp}}=\frac{E_{lp}-E_{gp}}{E_{gp}}$ (6)
where $E_{gp}$ denotes the energy of the global passive state. This normalized quantity offers a more physically consistent basis for comparison across states. In the following, we investigate whether this quantity can serve as the foundation for a thermodynamically motivated entanglement measure in the CV regime.
To this end we first derive an explicit expression for $\Delta\mathcal{E}_{\text{rel}}$ for a general two-mode Gaussian state. This can be achieved by employing the Bloch-Messiah decomposition [47], which parametrizes general two-mode Gaussian states in terms of elementary Gaussian operations implementable in quantum optical platforms. We consider such operations acting on an initial product state of two thermal modes with temperatures $T_{A}$ and $T_{B}$. The details of the decomposition, along with the derivation of the local and global passive energies, are presented in Appendices C and D. The resulting expression for the REG, as a function of the decomposition parameters, $\Delta\mathcal{E}_{\text{rel}}=f(k,\gamma,\alpha,\theta,z_{1},z_{2})$, is explicitly presented in Appendix D. Here, $k=\frac{k_{A}+k_{B}}{2}$ is the mean thermal fluctuation factor, $\gamma=\frac{k_{A}-k_{B}}{2}$ stands for the fluctuation difference, $\alpha=\frac{\omega_{B}}{\omega_{A}}$ is the ratio between the oscillation frequencies of the two modes, $z_{A}$ and $z_{B}$ are the respective squeezing parameters, and $\theta$ represents the beam splitter angle applied to entangle the modes.
Having introduced the REG, we now reveal a key structural distinction with respect to mutual information:
Lemma 2.
The relative ergotropic gap $\Delta\mathcal{E}_{\text{rel}}$ is functionally independent of the mutual information.
This independence, demonstrated in Appendix F, highlights the intrinsic significance of the REG and indicates that $\Delta\mathcal{E}_{\text{rel}}$ captures distinct features of quantum correlations not accounted for by mutual information alone. Motivated by this, we explore whether $\Delta\mathcal{E}_{\text{rel}}$ can discriminate between qualitatively distinct types of correlations, such as those in separable versus entangled states.
Theorem 2.
Let $\rho_{G}$ be a Gaussian state of two modes.
(a) If $\rho_{G}$ is separable, then $\Delta\mathcal{E}_{\text{rel}}(\rho_{G})$ is upper-bounded by
$\Delta\mathcal{E}_{\text{rel}}(\rho_{G})\leq\mathcal{B}^{\text{sep}}_{\mathrm{max}}(k,\gamma,\alpha)$ (7)
(b) If $\rho_{G}$ is entangled, then $\Delta\mathcal{E}_{\text{rel}}(\rho_{G})$ is lower-bounded by
$\Delta\mathcal{E}_{\text{rel}}(\rho_{G})>\mathcal{B}^{\text{ent}}_{\mathrm{min}}(k,\gamma,\alpha)$ (8)
The proof of the above theorem and explicit expressions for the bounds, together with a numerical illustration of the applicability of this result (see Fig.3) are provided in Appendix E. The two bounds are derived analytically employing the Positive Partial Transpose (PPT) criterion [48] as a necessary and sufficient condition for separability in Gaussian two-mode systems. Notably, they depend exclusively on the mean fluctuation parameters of the modes, their fluctuation difference, and the ratio of their respective frequencies. Recall that Lemma 1 establishes $\Delta\mathcal{E}_{\text{rel}}$ as strictly positive for all correlated states — whether classical or quantum — whereas Theorem 2 provides a sharper distinction via two analytic bounds: Eq.7 as a witness for entanglement, and Eq.8 as a witness for separability. These bounds satisfy $\mathcal{B}^{\text{ent}}_{\mathrm{min}}(k,\gamma,\alpha)\leq\mathcal{B}^{\text{sep}}_{\mathrm{max}}(k,\gamma,\alpha)$, but they do not always coincide. As a result, $\Delta\mathcal{E}_{\text{rel}}$ is a useful qualifier but not a full entanglement quantifier [49], since states in the ‘coexistence region’ [50, 32] characterized by values of $\Delta\mathcal{E}_{\text{rel}}$ between the two bounds cannot be unequivocally classified. However, the bounds do coincide in a physically relevant subclass of states.
Corollary 1.
Let $\rho_{G}$ be a Gaussian two-mode state such that the same fluctuation factor characterizes both modes. Then
$\Delta\mathcal{E}_{\text{rel}}(\rho_{G})\leq\mathcal{B}^{\text{sep}}_{\mathrm{max}}$ (9)
is a necessary and sufficient condition for separability.
The proof of this corollary is provided in Appendix E. As a consequence of these results, the faithfulness of $\Delta\mathcal{E}_{\text{rel}}$ as an entanglement measure is determined by the distance between the bounds $\mathcal{B}^{\text{sep}}_{\mathrm{max}}$ and $\mathcal{B}^{\text{ent}}_{\mathrm{min}}$: the closer they are, the more accurately our criterion distinguishes between separable and entangled states. In the limiting case where the two bounds coincide, the condition based on the REG becomes equivalent to the PPT criterion. To illustrate this behavior, Fig.1 shows the degree to which $\Delta\mathcal{E}_{\text{rel}}$ violates the separability bound for the parametric family of two-mode squeezed Gaussian states TMS($k,z$), for different values of the fluctuation difference $\gamma$.
Figure 1: Difference between $\mathcal{B}^{\text{sep}}_{\mathrm{max}}$ and $\Delta\mathcal{E}_{\text{rel}}$ for the parametric family of Gaussian two-mode squeezed states, as a function of the squeezing parameter $z$ and the mean fluctuations factor $k$. Values of fluctuation difference $\gamma$ have been set to {1,0.5,0} (left to right). The dashed and solid red lines represent the boundaries between separable and entangled states according to our criterion and the PPT condition, respectively. As predicted by Corollary 1, these boundaries coincide when $\gamma=0$.
Non-Gaussian states.– Gaussian states are experimentally and theoretically important, but their ease of characterization belies fundamental limitations in their practical usefulness for some quantum information tasks [51]. This is what motivates our interest in exploring the non-Gaussian realm [35] and to answer the question of how far the relationship between the REG and (quantum) correlations can be extended into it. By default, the results above (Theorems 1 and 2) directly apply to the class of non-Gaussian states that can be produced through local unitary operations on Gaussian ones. Outside this class, we examine the example of one photon-subtracted TMS states, whose REG can be computed from Eqs.3,5, and 6, exploiting the form of the covariance matrix derived in [52] (detailed derivations can be found in Appendix D). In Figure 2, $\Delta\mathcal{E}_{\text{rel}}$ is represented as a function of the same squeezing and fluctuation parameters $z,k$, respectively, of the initial TMS Gaussian state prior to the photon subtraction. Regarding the separability condition (dashed line in the figure), in the non-Gaussian case we use the criterion given by Shchukin and Vogel [36] as a benchmark, whose negativity is sufficient for entanglement. In particular, from the infinite set of inequalities given by such criterion, we pick the following:
$\langle\hat{a}^{\dagger}\hat{a}\rangle_{\rho}\langle\hat{b}^{\dagger}\hat{b}\rangle_{\rho}-\langle\hat{a}\hat{b}\rangle_{\rho}\langle\hat{a}^{\dagger}\hat{b}^{\dagger}\rangle_{\rho}<0\Longrightarrow\ \rho\ \text{entangled}$ (10)
where $\hat{a},\hat{b}$ are the annihilation operators associated to modes $A$ and $B$, respectively.
We observe numerically that there exists some trend by which entangled states (on the left of the separability boundary) display higher values of $\Delta\mathcal{E}_{\text{rel}}$, which increase the further they lie from the criterion’s dashed line. This behavior could allow us to identify the highest value of $\Delta\mathcal{E}_{\text{rel}}$ within the separability regime and establish it as an entanglement witness. We thus conjecture that this link between REG and entanglement observed in TMS photon-subtracted states could potentially be extended to other different classes of non-Gaussian states. The verification of this claim is left for future work.
Figure 2: REG of photon-subtracted TMS($z,k$) as a function of the fluctuation parameter $k$ of the initial thermal state and the parameter $z$ of the Gaussian squeezing transformation. The dashed line indicates the region of entangled states witnessed by the Shchukin-Vogel (SV) condition.
Discussions.– In this work we characterize correlations in CV systems via fundamental thermodynamic processes. Our results show that for bipartite Gaussian states, any correlation, even between asymmetric modes, ensures a strictly positive ergotropic gap, revealing a distinct form of correlation independent of entropy or informational quantities. Physically, this means that energy can be locked into correlations, making it inaccessible via local operations alone. For pure states, this phenomenon yields a valid entanglement quantifier, equivalent to the entropy of entanglement. However, for mixed states, the ergotropic gap diverges due to the unbounded energy spectrum of CV systems. To resolve this, we introduce the relative ergotropic gap (REG), which eliminates the divergence while preserving its role as a correlation indicator. This shows that entanglement shapes information and energy processing in a very different way for discrete and continuous systems. Notably, REG is functionally independent of mutual information, uncovering a distinct yet operationally meaningful aspect of quantum correlations. Furthermore, it enables a sharper classification of states, distinguishing entangled from separable ones through analytically derived bounds. While REG is not a universal entanglement measure, it provides a necessary and sufficient criterion under symmetric thermal conditions, paralleling the PPT test. When energy measurements are available [53], this could provide a more experimentally accessible entanglement criterion than full state tomography. Moreover, our results suggest REG’s broader applicability outside the Gaussian regime — an open direction for future study. Overall, our work establishes REG as an energy-based, experimentally accessible witness of quantum correlations, bridging thermodynamics and information-theoretic perspectives, and offering a novel tool for entanglement detection and characterization in near-term quantum optical technologies.
Acknowledgements.
Acknowledgements: This project has been funded by the Government of Spain (Severo Ochoa CEX2019-000910-S and FUNQIP), Fundació Cellex, Fundació Mir-Puig, Generalitat de Catalunya (CERCA program). F.C. acknowledges funding from the European Union (EQC, 101149233). M.A. acknowledges the funding supported by the European Union (Project- QURES, Grant Agreement No. 101153001). G.A. acknowledges funding from the UK Research and Innovation (UKRI EPSRC Grant No. EP/X010929/1). The views and opinions expressed in this Letter are, however, the sole responsibility of the author(s).
I Appendix A. Separability of Gaussian two-mode states
In the study of quantum information, the separability problem is a foundational question that addresses whether a given quantum state can be written as a convex mixture of product states. In this regard, the positive partial transpose (PPT) criterion [54, 55] is known to be a necessary condition for the separability of any bipartite quantum state, regardless of the Hilbert space dimension. In infinite-dimensional systems, the PPT condition can be evaluated on a bipartite state $\rho$ through its covariance matrix $\sigma$, yielding the following CV version of the criterion:
$\det(\sigma)-\det(\sigma_{A})-\det(\sigma_{B})+2\det(\sigma_{AB})+1\geq 0$ (11)
where $\sigma_{A}$ and $\sigma_{B}$ are the local covariance matrices of the two subsystems $A$ and $B$ of the bipartition, and $\sigma_{AB}$ represents their correlations:
$\sigma=\begin{pmatrix}\sigma_{A}&\sigma_{AB}\\ \sigma_{AB}&\sigma_{B}\end{pmatrix}$
In the context of Gaussian states, particularly in the two-mode case, it has been shown [48] that this condition is not only necessary but also sufficient for separability. Accordingly, if $\rho$ is a Gaussian state of two modes, and $\sigma_{sf}$ is the standard form of its covariance matrix:
$\sigma_{sf}=\begin{pmatrix}a&0&c_{1}&0\\ 0&a&0&c_{2}\\ c_{1}&0&b&0\\ 0&c_{2}&0&b\end{pmatrix},\quad\sigma_{A}=a\mathbb{1}_{2},\quad\sigma_{B}=b\mathbb{1}_{2},\quad\sigma_{AB}=\begin{pmatrix}c_{1}&0\\ 0&c_{2}\end{pmatrix},$
then the inequality
$\det(\sigma_{sf})-a^{2}-b^{2}+2c_{1}c_{2}+1\geq 0$ (12)
is satisfied if and only if $\rho$ is separable. Otherwise, it is entangled.
The applicability of this separability criterion extends beyond two-mode systems. In particular, it remains valid in the case of $1$ vs. $n$ mode bipartitions of Gaussian states [48], as well as for bisymmetric $(n+m)$ Gaussian states [32] (i.e., states that are invariant under mode permutations within any one of the two local subsystems). These generalizations provide a powerful framework for understanding entanglement structure in high-dimensional CV quantum systems while preserving computational tractability.
II Appendix B. Proof of Theorem 1
The covariance matrix of a Gaussian two-mode pure state is characterized by a single parameter, and in standard form it reduces to Eq.4 taking $a=b$, $c_{1}=-c_{2}=\sqrt{a^{2}-1}$:
$\sigma_{\text{pure}}=\left[\begin{matrix}a&0&\sqrt{a^{2}-1}&0\\ 0&a&0&-\sqrt{a^{2}-1}\\ \sqrt{a^{2}-1}&0&a&0\\ 0&-\sqrt{a^{2}-1}&0&a\end{matrix}\right],$ (13)
whereas through symplectic decomposition (i.e. diagonalization of $i\Omega\sigma_{\text{pure}}$) we obtain the eigenvalues $\nu_{+}=\nu_{-}=1$. Consequently, the ergotropic gap for these states reads
$\Delta\mathcal{E}_{\text{pure}}=\frac{\left(a-1\right)\left(\omega_{A}+\omega_{B}\right)}{2}.$ (14)
Additionally, from Eq.12 one can straightforwardly derive the separability condition for these states:
$\rho_{\text{pure}}\ \text{is separable}\iff 1\geq a^{2}$ (15)
Since by definition $1\leq a$, the only possibility for the RHS of Eq.15 to be fulfilled is $a=1$, and thus the only 2-mode Gaussian pure separable (i.e product) states are those whose local passive state is the vacuum (and the standard form of their covariance matrix is the identity). From here and the strict possitivity of the frequencies $\omega_{A},\omega_{B}$ it immediately follows that separability is a necessary and sufficient condition for $\Delta\mathcal{E}_{\text{pure}}=0$.
Lastly, for the ergotropic gap to be a consistent entanglement quantifier, it must be proven to be monotonic under $LOCC$. It turns out this can be shown along with the functional dependence between $\Delta\mathcal{E}_{\text{pure}}$ and the mutual information $I(A:B)$ between the two modes:
$I(A:B)=(a+1)log_{2}\bigg{(}\frac{a+1}{2}\bigg{)}-(a-1)log_{2}\bigg{(}\frac{a-1}{2}\bigg{)}$ (16)
which is known to be an entanglement monotone for pure states [56]. Indeed, there exists a monotonically increasing function $f$ that uniquely maps $\Delta\mathcal{E}_{\text{pure}}\ \rightarrow\ I(A:B)$, namely
$f(x)=(\gamma x+2)log_{2}\bigg{(}\frac{\gamma x+2}{2}\bigg{)}-(\gamma x)log_{2}\bigg{(}\frac{\gamma x}{2}\bigg{)}$ (17)
where we have taken $\gamma\equiv\frac{2}{\omega_{a}+\omega_{b}}$, and hence both metrics are LOCC monotones and provide equivalent information on the correlations of any pure two-mode Gaussian state. ∎
III Appendix C. Parametrization of Gaussian two-mode mixed states
Any Gaussian state $\rho$ can be fully described (up to displacements, which do not affect the ergotropic gap and are thus ignored here) by its covariance matrix (CM). In this appendix, we describe the so-called Bloch-Messiah decomposition [47, 29] of the most general CM of a two-mode Gaussian state. This decomposition provides a physically insightful way of expressing Gaussian states in terms of fundamental operations: local phase shifts, local squeezings, and two-mode beam splitter transformations. In this decomposition, any given covariance matrix $\sigma$ can be written as:
$\sigma=P(\varphi_{A},\varphi_{B})\ B(\theta)\ S(\sqrt{z_{A}},\sqrt{z_{B}})\ V(k_{A},k_{B})\ S^{T}(\sqrt{z_{A}},\sqrt{z_{B}})\ B^{T}(\theta)\ P^{T}(\varphi_{A},\varphi_{B})$ (18)
In what follows, we will adopt the $xpxp$ convention for mode ordering for the definition of the matrices appearing in the previous expression; that is, we will take the quadrature vector as $\hat{\vec{r}}=(\hat{x}_{A},\hat{p}_{A},\hat{x}_{B},\hat{p}_{B})^{T}$ as opposed to the alternative ordering $\hat{\vec{r}}=(\hat{x}_{A}\hat{x}_{B},\hat{p}_{A},\hat{p}_{B})^{T}$. The matrix
$V(k_{A},k_{B})=\begin{bmatrix}k_{A}&0&0&0\\ 0&k_{A}&0&0\\ 0&0&k_{B}&0\\ 0&0&0&k_{B}\end{bmatrix}\quad\quad\text{where}\ \ k_{A}\geq k_{B}\ \ \text{for}\ \ \omega_{A}\leq\omega_{B}$ (19)
is a diagonal matrix representing local thermal fluctuations. It encodes the symplectic eigenvalues of the state, which correspond to the effective thermal excitation levels of each mode. Specifically, $k_{A}$ and $k_{B}$ denote the thermal fluctuation parameters of modes A and B, respectively. $V(k_{A},k_{B})$ is itself the covariance matrix of a product of two uncorrelated thermal states, with each symplectic eigenvalue $k_{i}$ ($i\in{A,B}$) related to the corresponding temperature $T_{i}$ and mode frequency $\omega_{i}$ by
$k_{i}=\coth\bigg{(}\frac{\omega_{i}}{2K_{B}T_{i}}\bigg{)}\quad i\in\{A,B\}$ (20)
where $K_{B}$ denotes the Boltzmann constant.
For convenience, we define the mean and difference of the thermal fluctuations as
$k\equiv\frac{k_{A}+k_{B}}{2},\qquad\gamma\equiv\frac{k_{A}-k_{B}}{2}$ (21)
Here, $k$ quantifies the average thermal noise (mean excitation) present in the system, while $\gamma$ captures the asymmetry between the modes.
Matrix $S(\sqrt{z_{A}},\sqrt{z_{B}})$ represents local squeezing operations:
$S(\sqrt{z_{A}},\sqrt{z_{B}})=\begin{bmatrix}\sqrt{z_{A}}&0&0&0\\ 0&1/\sqrt{z_{A}}&0&0\\ 0&0&\sqrt{z_{B}}&0\\ 0&0&0&1/\sqrt{z_{B}}\end{bmatrix}$ (22)
with $\sqrt{z_{A}}$ and $\sqrt{z_{B}}$ the squeezing parameters for modes $A$ and $B$, respectively. These parameters are defined in terms of the effective quadrature rescaling due to local unitary operations (such as those generated by a single-mode squeezer).
In this parametrization, beam splitting is responsible for entanglement generation. The matrix associated to this transformation is:
$B(\theta)=\begin{bmatrix}\cos\theta&0&\sin\theta&0\\ 0&\cos\theta&0&\sin\theta\\ -\sin\theta&0&\cos\theta&0\\ 0&-\sin\theta&0&\cos\theta\end{bmatrix}$ (23)
which codifies the action of a beam splitter mixing the two modes, where $\theta\in[0,\pi/2]$ denotes the mixing angle, controlling the strength of mode entanglement. A value of $\theta=0$ corresponds to no interaction, while $\theta=\pi/4$ yields maximal mixing.
Lastly, $P(\varphi_{A},\varphi_{B})$ represents local phase rotations acting on each mode:
$P(\varphi_{A},\varphi_{B})=\begin{bmatrix}\cos\varphi_{A}&\sin\varphi_{A}&0&0\\ -\sin\varphi_{A}&\cos\varphi_{A}&0&0\\ 0&0&\cos\varphi_{B}&\sin\varphi_{B}\\ 0&0&-\sin\varphi_{B}&\cos\varphi_{B}\end{bmatrix}$ (24)
where $\varphi_{A}$ and $\varphi_{B}$ are the local phase shift parameters acting on modes $A$ and $B$, respectively. Since energy is invariant under local phase shifts, matrix $P(\varphi_{A},\varphi_{B})$ does not play a role in ergotropy-related considerations and its parameters do not influence our results.
The symplectic structure of the covaraince matrix does not depend on the physical frequencies of the modes. However, these frequencies enter the problem through the Hamiltonian and are necessary to define local energies and evaluate the ergotropic gap. Accordingly, we define:
$\omega\equiv\omega_{A}$: the natural oscillating frequency of mode $A$
$\alpha\equiv\frac{\omega_{B}}{\omega_{A}}$: the frequency ratio between the two modes, where $\alpha\geq 1$ since $\omega_{A}\leq\omega_{B}$.
While this appendix uses the Bloch-Messiah decomposition to express the Gaussian covariance matrix due to its physical transparency and modular structure, it is important to note that the results derived in the main text — particularly those concerning ergotropic gap bounds — are parametrization-independent. Any symplectic parametrization that captures the full Gaussian degrees of freedom would yield the same quantitative results. However, using the Bloch-Messiah form allows for a clear physical interpretation of each operation: local preparation (squeezing and thermal noise), mode interaction (beam splitting), and passive transformations (phase shifts).
IV APPENDIX D. Calculation of relative ergotropic gaps
Following Eq.6, the derivation of the relative ergotropic gap for a generic state $\rho$ reduces to finding its local and global passive energies. Moreover, since these energies depend only on first and second statistical moments (and first moments can be disregarded since they may be brought to zero through local Gaussian unitaries on the independent modes), the calculation of $\Delta\mathcal{E}_{\text{rel}}$ for any (not necessarily Gaussian) state depends solely on the entries of its covariance matrix.
IV.1 Gaussian two-mode states
Let $\sigma$ be the CM of a two-mode Gaussian state $\rho_{G}$, parametrized as explained in the previous Appendix. Then its corresponding global passive state is trivially that whose covariance matrix is
$\sigma_{\text{gp}}=\begin{bmatrix}k_{A}&0&0&0\\ 0&k_{A}&0&0\\ 0&0&k_{B}&0\\ 0&0&0&k_{B}\end{bmatrix}$ (25)
since this is the diagonal matrix of symplectic eigenvalues of $\sigma$. The global unitary that takes us from $\rho_{G}$ to its global passive is simply the inverse of the Gaussian transformations that define the parametrization of $\sigma$ (see Eq.18)
$\sigma_{\text{gp}}=G\ \sigma G^{T}$ (26)
where
$G=\bigg{(}P(\varphi_{A},\varphi_{B})\ B(\theta)\ S(\sqrt{z_{A}},\sqrt{z_{B}})\bigg{)}^{-1}$ (27)
and $\sigma$ is defined as in Eq.18. The global passive energy can be straightforwardly computed from the symplectic eigenvalues and Eq.3:
$E_{gp}=\frac{\omega}{2}[k_{A}+\alpha k_{B}-(1+\alpha)]=\frac{\omega}{2}[(k-1)(1+\alpha)+\gamma(1-\alpha)]$ (28)
For the local passive energy, we recall that any state whose covariance matrix is in standard form is locally passive. Therefore, our job reduces to finding the local Gaussian (symplectic) transformation that transforms $\sigma$ into $\sigma_{sf}$. The most general local Gaussian transformation can be decomposed in the product of single-mode squeezing and local rotations as:
$G_{loc}=P(\eta_{A},\eta_{B})\ S(r_{A},r_{B})\ P(\beta_{A},\beta_{B})$ (29)
When applying these local transformation $G_{loc}\ \sigma\ G_{loc}^{T}$, we obtain the product of two local phase shifts together, $P(\beta_{A},\beta_{B})$ and $P(\varphi_{A},\varphi_{B})$, which can be absorbed into a single matrix (to avoid extra degrees of freedom), or even chosen so that $\beta_{A}=-\varphi_{A},\beta_{B}=-\varphi_{B}$, since they are passive optics transformations that will not affect the energy. Same happens with the final phase shift $P(\eta_{A},\eta_{B})$, that will leave unchanged the energy of the state it is applied to, so we can also assume $\eta_{A}=0,\eta_{B}=0$. Finally, the standard form of $\sigma$ we are looking for takes the form:
$\sigma_{sf}=S(r_{A},r_{B})\ \sigma\ S(r_{A},r_{B})^{T}$ (30)
with $r_{A}$, $r_{B}$ taking values such that the resulting product is effectively in standard form. After some straight-forward algebra, one finds such form is only attained when
$r_{A}=\bigg{(}\frac{k_{A}z_{B}cos^{2}\theta+k_{B}z_{A}sin^{2}\theta}{(z_{A}z_{B})(k_{A}z_{A}cos^{2}\theta+k_{B}z_{B}sin^{2}\theta)}\bigg{)}^{1/4}$ (31)
and
$r_{B}=\bigg{(}\frac{k_{B}z_{A}cos^{2}\theta+k_{A}z_{B}sin^{2}\theta}{(z_{A}z_{B})(k_{B}z_{B}cos^{2}\theta+k_{A}z_{A}sin^{2}\theta)}\bigg{)}^{1/4}$ (32)
Substituting expressions 31,32 and 21 into Eq.30, we derive the explicit covariance matrix of the local passive state, whose energy, computed through Eq.3, yields:
E_lp = ω2 [ (k+γ)^2 cos^4θ+(k-γ)^2 sin^4θ+(k^2-γ^2)cos^2θsin^2θ(z12+ z22z1z2) + α(k-γ)^2 cos^4θ+(k+γ)^2 sin^4θ+(k^2-γ^2)cos^2θsin^2θ(z12+ z22z1z2) - (1+α) ]
and combining the local and global passive energies, the analytical expression for the relative ergotropic gap of a general two-mode Gaussian state yields
[…]
The Proof of Lemma 1 follows from the previous expression, where by inspection we conclude that the numerator of $\Delta\mathcal{E}_{\text{rel}}$, namely $\Delta\mathcal{E}$, vanishes is and only if $\theta=0$. In this case, there exist no correlations between the two modes of $\rho_{G}$, which can be written as a tensor product of the two. For any $\theta\in(0,2\pi)$, i.e., in the presence of (not necessarily quantum) correlations, the ergotropic gap is non-vanishing $\Delta\mathcal{E}>0$. ∎
IV.2 Non-Gaussian example: photon subtracted TMS state
The computation of the Gaussian relative ergotropic gap for a non-Gaussian state is exactly analogous to that of a Gaussian state, and again requires knowledge of the state’s covariance matrix exclusively. Such CM is not accessible in general, but it is known for some specific classes of non-Gaussian states. As an example, we focus on two-mode squeezed (TMS) states that undergo a photon subtraction (non-Gaussian operation) on their first mode. The covariance matrix $\sigma$ of the resulting state $\rho$ is derived in [52]:
$\sigma=\sigma_{0}+2\frac{(\sigma_{0}-\mathds{1})P(\sigma_{0}-\mathds{1})}{\text{Tr}[(\sigma_{0}-\mathds{1})P]}$ (35)
where $\sigma_{0}$ is the initial Gaussian TMS state’s covariance matrix and $P$ is the projection onto the subspace of the mode where the subtraction is performed. In our case (deterministic projection onto mode $A$):
$P=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix}$
The resulting $\sigma$ is in standard form (see Eq.4), with:
$a=k\bigg{(}z+\frac{1}{z}\bigg{)}-1,\quad\quad b=\frac{k\bigg{(}z^{2}+\frac{1}{z^{2}}\bigg{)}-\bigg{(}z+\frac{1}{z}\bigg{)}}{z+\frac{1}{z}-\frac{2}{k}},\quad\quad c_{1}=-c_{2}=k\bigg{(}-z+\frac{1}{z}\bigg{)}$
and its symplectic eigenvalues can be computed through
$\nu_{\pm}^{2}=\frac{\Gamma\pm\sqrt{\Gamma^{2}-4\det\sigma}}{2},\quad\Gamma=a^{2}+b^{2}+2c_{1}c_{2}$ (36)
Finally, the expression for the relative ergotropic gap is directly derived from the above values and Eqs.5 and 6.
V APPENDIX E. Proof of Theorem 2 and Corollary 1
Written in terms of the variables appearing in the Bloch-Messiah decomposition (Eq.18 with the substitution given by Eq.21), the separability condition in Eq. 12 for two-mode Gaussian states reads
$1+k^{4}-2k^{2}\gamma^{2}+\gamma^{4}-2k^{2}-2\gamma^{2}\geq 4\cos^{2}\theta\sin^{2}\theta\bigg{[}\left(k^{2}-\gamma^{2}\right)\left(\frac{z_{1}^{2}+z_{2}^{2}}{z_{1}z_{2}}\right)-2(k^{2}+\gamma^{2})\bigg{]}$ (37)
Part (a). First we note that by expansion of the binomial squares $(k+\gamma)^{2}$ and $(k-\gamma)^{2}$, the sum of square roots in the expression of the relative ergotropic gap (Eq.34) yields the following structure
$\sqrt{a+\delta}+\alpha\sqrt{a-\delta}\quad\text{with}\quad\delta=2k\gamma(\cos^{4}\theta-\sin^{4}\theta)$ (38)
Since $|\delta|\leq 2k\gamma$, we may construct the following chain of inequalities:
$\sqrt{a+\delta}+\alpha\sqrt{a-\delta}\leq(1+\alpha)\sqrt{a+|\delta|}\leq(1+\alpha)\sqrt{a+2k\gamma}$ (39)
and the relative ergotropic gap is upper bounded by
$\Delta\mathcal{E}_{\text{rel}}\leq\frac{-\Bigg{[}k(1+\alpha)+\gamma(1-\alpha)\Bigg{]}+(1+\alpha)\sqrt{(k^{2}+\gamma^{2})(\cos^{4}\theta+\sin^{4}\theta)+(k^{2}-\gamma^{2})\cos^{2}\theta\sin^{2}\theta\left(\frac{z_{2}^{2}+z_{1}^{2}}{z_{1}z_{2}}\right)+2k\gamma}}{(k-1)(1+\alpha)+\gamma(1-\alpha)}$ (40)
Since $\cos^{4}\theta+\sin^{4}\theta=1-2\cos^{2}\theta\sin^{2}\theta$:
$\Delta\mathcal{E}_{\text{rel}}\leq\frac{-\Bigg{[}k(1+\alpha)+\gamma(1-\alpha)\Bigg{]}+(1+\alpha)\sqrt{k^{2}+\gamma^{2}+2k\gamma+\cos^{2}\theta\sin^{2}\theta\bigg{[}(k^{2}-\gamma^{2})\left(\frac{z_{2}^{2}+z_{1}^{2}}{z_{1}z_{2}}\right)-2(k^{2}+\gamma^{2})\bigg{]}}}{(k-1)(1+\alpha)+\gamma(1-\alpha)}$ (41)
Finally, if we assume $\rho$ is separable, then we can embed the separability condition (Eq.37) into the previous inequality, and we are left with
$\Delta\mathcal{E}_{\text{rel}}^{\text{sep}}\leq\frac{-\Bigg{[}k(1+\alpha)+\gamma(1-\alpha)\Bigg{]}+\frac{1+\alpha}{2}\sqrt{1+k^{4}+\gamma^{4}+2k^{2}+2\gamma^{2}-2k^{2}\gamma^{2}+8k\gamma}}{(k-1)(1+\alpha)+\gamma(1-\alpha)}\equiv\mathcal{B}^{\text{sep}}_{\mathrm{max}}\quad\qquad\qed$ (42)
Part (b). Analogously to the previous proof, we proceed by merging the two square roots in Eq.34, through the inequality
$\sqrt{a+\delta}+\alpha\sqrt{a-\delta}\geq(1+\alpha)\sqrt{a-|\delta|}\geq(1+\alpha)\sqrt{a-2k\gamma}\qquad\text{since}\quad|\delta|=|2k\gamma(\cos^{4}\theta-\sin^{4}\theta)|\leq 2k\gamma$ (43)
with this and the trigonometry identity $\cos^{4}\theta+\sin^{4}\theta=1-2\cos^{2}\theta\sin^{2}\theta$, one arrives at:
$\Delta\mathcal{E}_{\text{rel}}\geq\frac{-\Bigg{[}k(1+\alpha)+\gamma(1-\alpha)\Bigg{]}+(1+\alpha)\sqrt{k^{2}+\gamma^{2}-2k\gamma+\cos^{2}\theta\sin^{2}\theta\bigg{[}(k^{2}-\gamma^{2})\left(\frac{z_{2}^{2}+z_{1}^{2}}{z_{1}z_{2}}\right)-2(k^{2}+\gamma^{2})\bigg{]}}}{(k-1)(1+\alpha)+\gamma(1-\alpha)}$ (44)
Since Eq.37 is an if and only if condition for separability, it can also be applied to characterize entangled states. If $\rho$ is entangled (not separable), then necessarily Eq.37 does not hold, and thus
$1+k^{4}-2k^{2}\gamma^{2}+\gamma^{4}-2k^{2}-2\gamma^{2}<4\cos^{2}\theta\sin^{2}\theta\bigg{[}\left(k^{2}-\gamma^{2}\right)\left(\frac{z_{1}^{2}+z_{2}^{2}}{z_{1}z_{2}}\right)-2(k^{2}+\gamma^{2})\bigg{]}$ (45)
Embedding this new inverse inequality into Eq.44, we finally obtain the lower bound on $\Delta\mathcal{E}_{\text{rel}}$ for entangled states:
$\Delta\mathcal{E}_{\text{rel}}^{\text{ent}}>\frac{-\Bigg{[}k(1+\alpha)+\gamma(1-\alpha)\Bigg{]}+\frac{1+\alpha}{2}\sqrt{1+k^{4}+\gamma^{4}+2k^{2}+2\gamma^{2}-2k^{2}\gamma^{2}-8k\gamma}}{(k-1)(1+\alpha)+\gamma(1-\alpha)}\equiv\mathcal{B}^{\text{ent}}_{\mathrm{min}}\quad\qquad\qed$ (46)
The following figure illustrates the meaning of these two bounds. Relative ergotropic gap is shown for $N=5\times 10^{2}$ randomly generated two mode Gaussian states. The blue dashed line represents the minimum $\Delta\mathcal{E}_{\text{rel}}$ for entangled states, while the upper bound for separable states is displayed in dashed red. Since the two bounds $\mathcal{B}^{\text{ent}}_{\mathrm{min}}$ and $\mathcal{B}^{\text{sep}}_{\mathrm{max}}$ are dependent on $k,\gamma$ and $\alpha$, the values of these parameters have been fixed so a single threshold holds for all of the simulated states. Neither entanglement nor separability of states with REG lying in between the two lines can be witnessed by means of our criterion. However, any value of REG above the red line or below the blue one can unambiguously certify the state is entangled or separable, respectively.
Figure 3: REG for $N=5\times 10^{2}$ random two-mode Gaussian states with fixed parameters $k=2.5$, $\gamma=0.5$ and $\alpha=10$. The blue and red dashed lines represent the lower and upper bounds for entangled and separable states, respectively.
Proof of Corollary 1:
If $\rho_{G}$ has equal fluctuation factors $k_{A},k_{B}$ for both modes, then $k=k_{A}=k_{B}$ and $\gamma=0$. In that case, $\mathcal{B}^{\text{sep}}_{\mathrm{max}}=\mathcal{B}^{\text{ent}}_{\mathrm{min}}$. From part (a) of Theorem 2 we derive that Eq.9 is necessary for separability. Sufficiency follows from part (b). ∎
VI APPENDIX F. Independency of relative ergotropic gap
We now demonstrate that, in contrast to the pure-state scenario, the relative ergotropic gap remains a distinct and informative quantity for mixed states. In particular, it captures aspects of correlations not accounted for by the mutual information. To assess whether these two quantities are functionally dependent, we analyze their gradients with respect to the set of variables appearing in the standard form of a generic covariance matrix $\mathbf{x}=(a,b,\nu_{+},\nu_{-})$. Functional dependence between two scalar functions $f(\mathbf{x})$ and $g(\mathbf{x})$ can be tested by examining their Jacobian matrix:
$J=\begin{bmatrix}\frac{\partial f}{\partial a}&\frac{\partial f}{\partial b}&\frac{\partial f}{\partial\nu_{+}}&\frac{\partial f}{\partial\nu_{-}}\\ &\\ \frac{\partial g}{\partial a}&\frac{\partial g}{\partial b}&\frac{\partial g}{\partial\nu_{+}}&\frac{\partial g}{\partial\nu_{-}}\end{bmatrix}$ (47)
If this matrix has full row rank (i.e., rank 2) at a certain point, the functions $f$ and $g$ are functionally independent at that point in parameter space. We proceed by explicitly computing the gradients of both the relative ergotropic gap $\Delta\mathcal{E}_{\text{rel}}$ and the mutual information $I(A\!:\!B)$, and then evaluating the Jacobian matrix.
From Eqs.5 and 6, we derive the expression
$\Delta\mathcal{E}_{\text{rel}}=\frac{\omega_{A}(a-\nu_{+})+\omega_{B}(b-\nu_{-})}{\omega_{A}(\nu_{+}-1)+\omega_{B}(\nu_{-}-1)}$ (48)
we will assume $\omega_{A}=\omega_{B}$, since it does not affect the generality of the result. The corresponding gradient reads
$\nabla\bigg{[}\Delta\mathcal{E}_{\text{rel}}\bigg{]}=\bigg{(}\frac{\partial\Delta\mathcal{E}_{\text{rel}}}{\partial a},\frac{\partial\Delta\mathcal{E}_{\text{rel}}}{\partial b},\frac{\partial\Delta\mathcal{E}_{\text{rel}}}{\partial\nu_{+}},\frac{\partial\Delta\mathcal{E}_{\text{rel}}}{\partial\nu_{-}}\bigg{)}=\bigg{(}\frac{1}{\nu_{+}+\nu_{-}-2},\ \frac{1}{\nu_{+}+\nu_{-}-2},\ \frac{2-b-a}{(\nu_{+}+\nu_{-}-2)^{2}},\ \frac{2-b-a}{(\nu_{+}+\nu_{-}-2)^{2}}\bigg{)}$ (49)
and for the mutual information, defined as
$I(A:B)=\frac{\sum_{x\in\{a,b\}}(x+1)log_{2}(\frac{x+1}{2})-(x-1)log_{2}(\frac{x-1}{2})-\sum_{y\in\{\nu_{+},\nu_{-}\}}(y+1)log_{2}(\frac{y+1}{2})-(y-1)log_{2}(\frac{y-1}{2})}{2}$ (50)
the gradient takes the form
$\nabla I(A:B)=\bigg{(}\frac{log_{2}(\frac{a+1}{2})-log_{2}(\frac{a-1}{2})}{2},\frac{log_{2}(\frac{b+1}{2})-log_{2}(\frac{b-1}{2})}{2},\frac{log_{2}(\frac{\nu_{+}-1}{2})-log_{2}(\frac{\nu_{+}+1}{2})}{2},\frac{log_{2}(\frac{\nu_{-}-1}{2})-log_{2}(\frac{\nu_{-}+1}{2})}{2}\bigg{)}$ (51)
By inspection of both gradients, one can immediately conclude they are independent vectors (it suffices to check that the first two entries of $\nabla\bigg{(}\Delta\mathcal{E}_{\text{rel}}\bigg{)}$ have the same value, whereas for $a\neq b$,
$\frac{log_{2}(\frac{a+1}{2})-log_{2}(\frac{a-1}{2})}{2}\neq\frac{log_{2}(\frac{b+1}{2})-log_{2}(\frac{b-1}{2})}{2}$ (52)
Consequently, the Jacobian matrix (see Eq.47) has rank 2, and the functions are proven to represent independent quantifiers of entanglement correlations for mixed states. It is important to note that the consideration of the particular case $\omega_{A}=\omega_{B}$ and $a\neq b$ in the above proof preserves the validity of the result, since functional independency of two functions at one single point in parameter space suffices to prove independency in general.   ∎